OpenStudy (anonymous):

PLEASE HELP **MEDAL AND FAN** Solve, using the substitution method. 2x + 3y = 1 y = x – 8 Question 3 options: (0, –8) (2, –6) (4, –4) (5, –3)

OpenStudy (anonymous):

Choose one equation from each pair and solve for y. Then substitute that y into the second equation. IE 3x+y=4 x+2y=7 Solving for y the first equation becomes y=-3x+4. Plug that in for y in the second equation: x+2(-3x+4)=7 Simplify for x. -5x=1 x=-1/5 choose one of original equations and plug in x and solve for y. 3(-1/5)+y=4 y=23/5

OpenStudy (mathmale):

Post one question at a time, please. Have you used the substitution method before? If so, describe what you did and why.

OpenStudy (anonymous):

It is one question? @mathmale

OpenStudy (mathmale):

Then, ask yourself: "What would be a suitable substitution to use in Problem 1?"

OpenStudy (mathmale):

Yes, you're right; it is just one question. Sorry. :(

OpenStudy (anonymous):

Its fine!

OpenStudy (anonymous):

kylie i answered your question its at the top

OpenStudy (anonymous):

I didnt understand it so then whats the answer? @Cj_2COOL

OpenStudy (anonymous):

-x+3y=4 x+6y=14 Take an equation: x + 6y = 14 Then make it equal x. x + 6y = 14 - 6y -6y x = - 6y + 14 Then plug x into the first equation that was not used: -(- 6y + 14) + 3y = 4 6y - 14 + 3y = 4 9y - 14 = 4 +14 +14 9y = 18 /9 /9 y = 2 Then plug y into either equation -x+3y=4 -x + 3(2) = 4 -x + 6 = 4 -6 -6 -x = -2 /-1 /-1 x = 2 The solution for the first problem is x = 2 and y = 2

OpenStudy (mathmale):

Thanks for your input, @Cj_2COOL , but next time do NOT do the problem for the other person; instead, provide structure and guidance so that that person can find his or her own answer.

OpenStudy (anonymous):

Thank you for both of your help but i think i found the answer is it D?

OpenStudy (mathmale):

I'd be glad to respond if you'd share your work first. Need to know how you decided upon an answer.

OpenStudy (anonymous):

Sorry mathmale ill remember that

OpenStudy (mathmale):

:)

OpenStudy (anonymous):

I just thought of (x,y) replace the #'s in with x,y

OpenStudy (mathmale):

But you're supposed to use substitution specifically. You will be asked in the future to use this method. You are given 2 equations. Which one has already been solved for y?

OpenStudy (anonymous):

The second one

OpenStudy (mathmale):

right. That equation states that y is the same as (x-8). Throw out y in the first equation and replace it with (x-8). (You must keep those parentheses.) Show your results.

OpenStudy (mathmale):

\[2x+3y=1\rightarrow 2x + 3( ?? )=1\]

OpenStudy (anonymous):

Thank you for everything but its fine!

OpenStudy (mathmale):

Sorry, but I need to get off the 'Net. Bye.